What are the forces on two charges of +0.50 C and +1.7 C, respectively, if they are separated by a distance of 3.3 m?
_____ N

Do the charges attract or repel each other?
Answer #1

Apply

F = k*q1*q2/r^2

here q1, q2 are the magnitude of thecharges and r is the distance between them.

so,
F = 9*10^9*0.5*1.7/3.3^2

= 7.02*10^8 N

This force is repulsive.

Beacuse like charges repel and opposite charge attract.